# MMTU: A Massive Multi-Task Table Understanding and Reasoning Benchmark

### v2

@@ 4 Experiments
Experimental setup. In all our experiments reported below, we use publicly available model endpoints for inference aoa ; fou with default parameter settings. All of our code and data are publicly available at cod for future research.
@@ 4.1 Overall performance
Chat Qwen3-235B-A22B-Instruct-2507 0.509 ± 0.01 0.00044
GPT-4o (2024-11-20) 0.491 ± 0.01 0.01019
Llama-4-Maverick-17B-128E-Instruct-FP8 0.472 ± 0.01 0.00066
Llama-3.3-70B 0.438 ± 0.01 0.0015
Mistral-Large-2411 0.430 ± 0.01 0.01066
Mistral-Small-2503 0.402 ± 0.01 0.00273
GPT-4o-mini (2024-07-18) 0.386 ± 0.01 0.00061
Llama-4-Scout-17B-16E-Instruct 0.377 ± 0.01 0.00036
Qwen3-32B (no thinking) 0.367 ± 0.01 0.00007
Qwen3-8B (no thinking) 0.343 ± 0.01 0.00015
Qwen2.5-7B-Instruct 0.300 ± 0.01 0.00010
Llama-3.1-8B 0.259 ± 0.01 0.00003
Reasoning o3 0.671 ± 0.01 0.01539
Gemini-2.5-pro 0.646 ± 0.01 0.00790
o4-mini (2024-07-18) 0.639 ± 0.01 0.00993
Grok-3-mini 0.624 ± 0.01 0.00111
Gemini 2.5 Flash 0.606 ± 0.01 0.00199
Deepseek-R1 0.596 ± 0.01 0.00167
Grok-3 0.576 ± 0.01 0.01232
gpt-oss-120b 0.523 ± 0.01 0.00050
Qwen3-235B-A22B-Thinking-2507 0.510 ± 0.01 0.00068
Qwen3-32B (thinking) 0.488 ± 0.01 0.00017
gpt-oss-20b 0.460 ± 0.01 0.00040
Qwen3-8B (thinking) 0.460 ± 0.01 0.00041
@@
Table 3: Overall performance and cost results of chat and reasoning models. Cost per question are calculated based on public pricing information on https://openrouter.ai/ as of Aug 18, 2025.
@@
We benchmark a range of frontier open-source and proprietary models using MMTU. Table 3 gives an overview of their performance, as well as their cost information. Notably, the two reasoning-focused models, OpenAI o4-mini and DeepSeek R1, perform the best at 63.9% and 59.6% respectively, significantly outperforming the chat-based models and underscoring the challenging nature of MMTU. Our analysis of the reasoning traces generated by R1 reveals that reasoning models excel on MMTU because of their strong coding skills and their ability to break complex tasks over large tables, into a sequence of more manageable sub-tasks on smaller table context (i.e., subsets of rows and columns relevant to the task at hand).
@@
From a cost-efficiency perspective [^1] Price-per-token figures are sourced from the respective API providers pri (a, b, c, d, e, f, g, h)., both reasoning models are also competitive (despite producing more intermediate thinking tokens), due to the relatively light-weight nature of R1 and o4-mini, making them cost-effective choices for tackling table tasks.
@@
In the rest of this paper, we will focus on four best-performing frontier models: two reasoning models (DeepSeek R1 and OpenAI o4-mini), and two general-purpose models (GPT-4o and Llama 3.3 70B), for detailed analysis to avoid clutter.
@@
Figure 5 provides a detailed comparison across 10 task categories. While reasoning models outperform chat-based models, the relative empty nature of the radar chart reveals substantial gaps in model performance, highlighting opportunities for improvement. For instance, we can see that models still face difficulties with table-centric coding tasks (e.g., Column Transform, Table Transform, etc.). Tasks that require holistic reasoning across multiple tables and multiple columns (e.g., Table Join, Column Relationship, Data Cleaning, etc.) also remain challenging. We will present an error-analysis in Section 4.3.
@@
Figure 4: Performance comparison in all 25 tasks, for chat-models GPT-4o and Llama-3 (solid bars), as well as reasoning models DeepSeek and OpenAI o4-mini (shaded bars). Reasoning models are noticeably better on average.
@@
4.2 Detailed analysis and sensitivity
These findings suggest that, despite recent advances in long-context LLMs Chen et al. (2023); Achiam et al. (2023), long contexts remain a significant challenge for tables. In Appendix C, we use a detailed experimental comparison between the standard “needle-in-a-haystack (NIH)” NIA ; Roberts et al. (2024), and our table-based “needle-in-a-haystack in table (NIHT)” test – while frontier models perform nearly perfectly on NIH, their performance drops sharply on NIHT, revealing fundamental limitations in their ability to handle long table contexts.
Robustness to table permutation. Figure 7 illustrates model performance under different table permutations. Specifically, we randomly shuffle rows and/or columns in input tables [^2] In shuffling columns, we keep the 3 left-most columns in all tables intact, as these are likely the “key columns” or “entity columns” in tables Chakrabarti et al. (2016); Cafarella et al. (2008a) in tables, to best preserve the meaning of these tables., with the 4 bars in each group representing: (1) no shuffle, (2) row-only shuffle, (3) column-only shuffle, and (4) both row and column shuffle. Recall that unlike natural-language text, two dimension relational tables are permutation-invariant Li et al. (2023c); Cong et al. (2023), meaning that shuffling rows and columns should generally not change their semantic meanings and the associated tasks (e.g., the tasks in Figure 1 will remain the same, even if the rows/columns in the associated tables are shuffled).
Table format variations. Figure 8 shows different models’ performance using common table input formats: Markdown, CSV, JSON, and HTML. We observe that unlike prior studies that showed notable sensitivity of LLMs Sui et al. (2024) to table formats, our results indicate that today’s frontier models, especially reasoning ones (o4-mini and R1), are becoming less sensitive to these format variations, reducing the need for format-specific optimization as models continue to improve. For chat models (GPT-4o and Llama-3.3), we note that using JSON still has an advantage on MMTU, mainly because it is easier for models to identify value/column correspondence in the JSON format, especially in long table context settings (e.g., in needle-in-a-haystack tasks like NIHT and NIHI shown in Table 1).
4.3 Error analysis
We observe that models are prone to errors when handling long table contexts, such as multiple tables or large tables with many rows and columns, as illustrated in Figure 6. Figure 10 shows a concrete example of the issue in the task of table-reshaping (TTBR in Table 1), where the reasoning trace shows that the model miscalculates the column index of the target column. This issue of long-context table is also notable in tasks like List-to-Table (L2T), Data Imputation (DI), Equi-Join (EJ), and NL-to-SQL (NS), etc.
A more detailed analysis of challenges in long-context tables is provided in Appendix C, where we examine a table-specific variant of the “needle-in-a-haystack” task, to understand why long-context tables remain challenging for models.
Figure 9: An example “table understanding” error in Data Imputation: while the model knows the person’s name for the missing cell, it misses the table context and uses an abbreviated form of the name, instead of the full-name in the ground-truth.
Figure 10: An example “table understanding” error in Table Reshaping: while the model knows the correct transforms to perform (“explode” in Pandas), it miscounts the column index for the target column “Season(s)”, which should be the sixth column instead of the fifth based on the reasoning trace, which leads to an incorrect transformation.
5 Conclusion
Future work includes broadening the scope of table tasks to cover those that are important in practice but underrepresented in the existing research literature, as well as incorporating more subjective or creative tasks such as data generation, summarization, and enrichment. Another promising direction is the integration of multi-modal evaluation.
Appendix A Task overview
A.1 Column transformation
Program Transform by Example (PTBE) He et al. (2018); Gulwani (2011); Devlin et al. (2017); Singh (2016). This is a popular task studied in the data management and programming language literature. In this task, we are given a few example output values in a new output column (usually provided by users as demonstrations), and the task is to synthesize a transformation program using input tables, such that the produced output can match the user-given output examples. Figure 1 shows an example of this task. We use the benchmark datasets from He et al. (2018); Gulwani (2011) for this task in MMTU.
Semantic Transform by example (STBE) He et al. (2015); Abedjan et al. (2016); Dargahi Nobari and Rafiei (2024). This task is similar to PFBE above, in that users also provide a few example output to demonstrate the intended transformations, but the target transformation requires “semantic” transformations (e.g., country to capital-city, or company to stock-ticker), that cannot be programmed using a syntactic program like in PFBE. We use the datasets from He et al. (2015); Abedjan et al. (2016)for our benchmark.
Formula by Context (FBC) Chen et al. (2024a, 2021b). This task is studied in the context of spreadsheets, where given a target spreadsheet cell in which users want to enter a spreadsheet formula, we are asked to predict the intended formula in the target cell. We use the 4 benchmark datasets in Chen et al. (2024a).
A.2 Table Transformation
Table Transform by Output Table (TTBT) Bavishi et al. (2019); Wang et al. (2017); Tran et al. (2009); Zhang and Sun (2013). In this task, we are given a pair of input/output tables, and the task is to infer a transformation program (in SQL or Pandas), which when executed on the input table, can generate the desired output table. Figure 1 shows an example of this task. We use the datasets from Bavishi et al. (2019); Wang et al. (2017).
Table Transform by Output Schema (TTBS) Yang et al. (2021). This task is similar to TTBT above, except that the output table is a schematic depiction of how the desired output table should look like, without being the exact output that the desired transformation program would generate on the given input (as it is sometimes hard to generate such output table without first producing the desired program). We use the dataset in Yang et al. (2021)for this task.
Table Transform by Relationalization (TTBR) Li et al. (2023b); Barowy et al. (2015); Jin et al. (2017). Because relational data analysis often requires input tables to be in a relational form, this task transforms data in non-relational semi-structured forms, into the standard relational form. The task is to predict such a relationalization transformation program, based on the characteristics of the input table. We use the dataset in Li et al. (2023b).
A.3 Table matching
Entity Matching (EM) Das et al. ; Mudgal et al. (2018); Zhao and He (2019); Li et al. (2020); Papadakis et al. (2021). Entity matching is the task of determining whether rows or entities from two tables correspond to the same real-world entity. We use the datasets in  Das et al. ; Mudgal et al. (2018)for benchmarking.
Schema Matching (SM) Koutras et al. (2021); Zhang et al. (2024); Bellahsene et al. (2011); Madhavan et al. (2001). Schema matching tries to match relevant columns from two tables that map to the same semantic concept. Figure 1 shows an example of this task. We use datasets in Koutras et al. (2021); Zhang et al. (2024)for MMTU.
Header Value Matching (HVM) Li et al. (2023c); Rahm and Bernstein (2001). In HVM, we are given a table without headers, and a shuffled list of the original headers in the table, where the task is to match column headers with column values in tables. We use the dataset in Li et al. (2023c)for benchmarking.
A.4 Data cleaning
Data Imputation (DI) Cafarella et al. (2008a); Eggert et al. (2023); Li et al. (2023c). Data imputation is the intuitive task of predicting missing values in a relational table, based on the surrounding context of the table. Figure 1 shows an example of this task. We use the dataset from Cafarella et al. (2008a); Eggert et al. (2023).
Error Detection (ED) Chu et al. (2016); Wang and He (2019); Mahdavi and Abedjan (2020); Chen et al. (2025). The task of Error detection aims to identify erroneous values in a table that are semantically inconsistent or anomalous with the rest of the column. We use the dataset in Chen et al. (2025).
List to Tables (L2T) Chu et al. (2015); Cafarella et al. (2008b); Elmeleegy et al. (2009); Lerman et al. (2001). In List-to-table, the task is to segment records of data without clear separators, into columns, so that values in the same column are homogeneous, and the resulting table becomes a natural relational table, with values consistently aligned in homogenous columns. We use the dataset from Chu et al. (2015).
A.5 Table join
Equal Join (EJ) Lin et al. (2023); Chen et al. (2014); Rostin et al. (2009); Jiang and Naumann (2020). In Equal-join, we are given a collection of related tables, and the task is to identify all join relationships between the tables. Figure 1 shows an example of this task. We use the dataset in Lin et al. (2023).
Semantic Join (SJ) He et al. (2015); Abedjan et al. (2016); Dargahi Nobari and Rafiei (2024); Wang and He (2017). In Semantic Join, we are also tasked to join related tables together, but instead of the common equi-join, which is based on string-equality comparisons, the join relationship is based on semantic relatedness (e.g., country and capital city, or company and stock ticket). We use the same semantic-transformation datasets from He et al. (2015); Abedjan et al. (2016), but converting the input/output transformation columns, as join keys from two separate tables.
A.6 Column relationships
Arithmetic Relationship (AR) Shraga and Miller (2023); Aut . This task focuses on identifying arithmetic relationships from an underlying table. Figure 1 shows an example of this task.
String Relationship (SR) Aut ; He et al. (2018); Gulwani et al. (2012). This task focuses on identifying string transformation relationships from a given table.
Functional Relationship (FR) Aut ; Parciak et al. (2024); Wei and Link (2019). This task focuses on identifying functional-relationships from input tables. We use the datasets from Aut our as benchmarks for all three tasks.
A.7 Table understanding
Needle-in-a-haystack-table (NIHT). In this task, we create a variant of the popular “Needle-in-a-haystack” task in the context of tables, like described in detail in Appendix C. We use the dataset from Eggert et al. (2023)to construct tests in this task.
Needle-in-a-haystack-index (NIHI). In this task, we reverse the NIHT task above, and ask models to identify index positions corresponding to a given value in a table. We use the same dataset from Eggert et al. (2023)to construct tests for this task.
A.8 NL-2-Code
NL-2-SQL (NS) Zheng et al. (2024); Li et al. (2024); Lee et al. (2021); Yu et al. (2018); Zhong et al. (2017). NL-2-SQL is a popular task to translate natural-language questions into SQL statements that can execute given an input table. We use the benchmarks from Zheng et al. (2024); Li et al. (2024); Lee et al. (2021); Yu et al. (2018); Zhong et al. (2017)in MMTU.
A.9 Table Question Answering (Table QA)
Table-QA (TQA) Wu et al. (2025); Yang et al. (2015); Chen et al. (2021c); Cheng et al. (2021). Table QA is another popular task, where models are used to directly answer questions on a given input table, without code execution. We use benchmarks from Chen et al. (2019); Wu et al. (2025); Yang et al. (2015); Chen et al. (2021c)for this task.
Fact verification (FV) Chen et al. (2019); Yang and Zhu (2021); Zhang et al. (2020a). Table fact verification is a variant of TQA, in which models are asked if a statement is refuted or supported by facts presented in an input table. We use data from Chen et al. (2019)for this task.
A.10 Knowledge-base mapping (KB Mapping)
Column type annotation (CTA) Jiménez-Ruiz et al. (2020); Yan and He (2018); Suhara et al. (2022); Zhang et al. (2019); Hulsebos et al. (2019). This is a popular task where columns in a table is mapped to known entity types in a knowledge-base. We use the CTA dataset from Jiménez-Ruiz et al. (2020).
Column property annotation (CPA) Jiménez-Ruiz et al. (2020); Cutrona et al. (2022); Nguyen et al. (2021). Jiménez-Ruiz et al. (2020). The CPA task is to predict the relationship of a given pair of columns, based on known properties in a knowledge-base. We use the CPA dataset from Jiménez-Ruiz et al. (2020).
Cell entity annotation (CEA) Jiménez-Ruiz et al. (2020); Deng et al. (2022); Bhagavatula et al. (2015); Wu et al. (2016); Liu et al. (2023). The CEA task predicts the knowledge-base entity id of a given cell in a table. We use the CEA dataset from Jiménez-Ruiz et al. (2020).

### v3

@@ 4 Experiments
Experimental setup. In all our experiments reported below, we use publicly available model endpoints for inference [aoai-api, foundry-api] with default parameter settings. All of our code and data are publicly available at [code-data] for future research.
@@ 4.1 Overall performance
Reasoning GPT-5 0.696 ± 0.01 0.01727
o3 0.691 ± 0.01 0.01539
GPT-5-mini 0.667 ± 0.01 0.00276
Gemini-2.5-pro 0.665 ± 0.01 0.00790
o4-mini (2024-07-18) 0.660 ± 0.01 0.00993
Grok-3-mini 0.645 ± 0.01 0.00111
Gemini 2.5 Flash 0.625 ± 0.01 0.00199
Deepseek-R1 0.579 ± 0.01 0.00167
gpt-oss-120b 0.543 ± 0.01 0.00050
Qwen3-235B-A22B-Thinking-2507 0.529 ± 0.01 0.00068
Qwen3-32B (thinking) 0.506 ± 0.01 0.00017
gpt-oss-20b 0.478 ± 0.01 0.00040
Qwen3-8B (thinking) 0.473 ± 0.01 0.00041
Chat GPT-5-Chat 0.577 ± 0.01 0.00534
Deepseek-V3 0.555 ± 0.01 0.00095
Qwen3-235B-A22B-Instruct-2507 0.524 ± 0.01 0.00044
GPT-4o (2024-11-20) 0.507 ± 0.01 0.01019
Llama-4-Maverick-17B-128E-Instruct-FP8 0.490 ± 0.01 0.00066
Llama-3.3-70B 0.454 ± 0.01 0.0015
Mistral-Large-2411 0.446 ± 0.01 0.01066
Mistral-Small-2503 0.417 ± 0.01 0.00273
GPT-4o-mini (2024-07-18) 0.400 ± 0.01 0.00061
Llama-4-Scout-17B-16E-Instruct 0.393 ± 0.01 0.00036
Qwen3-32B (no thinking) 0.379 ± 0.01 0.00007
Qwen3-8B (no thinking) 0.353 ± 0.01 0.00015
Qwen2.5-7B-Instruct 0.310 ± 0.01 0.00010
Llama-3.1-8B 0.268 ± 0.01 0.00003
@@
Table 3: Overall performance and cost results of chat and reasoning models. Cost per question is calculated based on public pricing information on https://openrouter.ai/ as of Aug 2025. Results for top-performing models, such as GPT-5 and GPT-5-Chat, are averaged over 3 runs.
We benchmark a range of frontier open-source and proprietary models using MMTU. Table 3 gives an overview of their performance, as well as their cost information. Notably, reasoning models such as GPT-5, o3 and Gemini-2.5-pro, significantly outperform chat-oriented models, with OpenAI GPT-5 achieving the best score at 69.6%, highlighting the challenging nature of MMTU. Among open-source reasoning models, DeepSeek R1 attains the best performance with a score of 57.9%, which while strong, reveals a potential gap between leading proprietary and open-source models.
@@
Our analysis of traces generated by reasoning models reveals that they excel on MMTU over chat models, because of their strong coding skills, and abilities to break complex tasks on large tables, into a sequence of more manageable sub-tasks with smaller table context (i.e., subsets of rows and columns relevant to the task at hand).
@@
From a cost-efficiency perspective [^1] Price-per-token figures are sourced from the respective API providers [price-deepseek-r1, price-gpt-4o, price-gpt-4o-mini, price-llama3.1-8B, price-llama3.3-70B, price-mistral-large-2411, price-mistral-small-2503, price-o4-mini]., looking at the “cost per question” column in the table, several light-weight reasoning models (e.g., GPT-5-mini and Gemini-2.5-flash) are quite competitive in terms of both quality and cost (even after accounting for their intermediate thinking tokens), making them cost-effective choices for complex table tasks.
@@
While we benchmark a broad range of models, our detailed analysis in the remainder of the paper will focus on four leading models to avoid clutter: two reasoning models (OpenAI GPT-5 and DeepSeek R1) and two chat-oriented models (OpenAI GPT-5-Chat and DeepSeek V3).
@@
Figure 5 provides a detailed comparison across 10 task categories, between these four leading models. While reasoning models outperform chat-based models, the relative empty nature of the radar chart reveals substantial gaps in model performance, highlighting opportunities for improvement. For instance, we can see that models still face difficulties with table-centric coding tasks (e.g., Column Transform, Table Transform, etc.). Tasks that require holistic reasoning across multiple tables and multiple columns (e.g., Table Join, Column Relationship, Data Cleaning, etc.) also remain challenging. We will present an error-analysis in Section 4.3.
@@
Figure 4: Performance comparison in all 25 tasks, for chat-models GPT-5-Chat and DeepSeek-V3 (solid bars), as well as reasoning models DeepSeek-R1 and OpenAI GPT-5 (shaded bars). On average, reasoning models are noticeably better.
@@
4.2 Detailed analysis and sensitivity
These findings suggest that, despite recent advances in long-context LLMs [long-context-2, Gpt-4-tr], long contexts remain a significant challenge for tables. In Appendix E, we use a detailed experimental comparison between the standard “needle-in-a-haystack (NIH)” [NIAH-orig, niah-2], and our table-based “needle-in-a-haystack in table (NIHT)” test – while frontier models perform nearly perfectly on NIH, their performance drops sharply on NIHT, revealing fundamental limitations in their ability to handle long table contexts.
Robustness to table permutation. Figure 7 illustrates model performance under different table permutations. Specifically, we randomly shuffle rows and/or columns in input tables [^2] In shuffling columns, we keep the 3 left-most columns in all tables intact, as these are likely the “key columns” or “entity columns” in tables [bing-assets, webtables] in tables, to best preserve the meaning of these tables., with the 4 bars in each group representing: (1) no shuffle, (2) row-only shuffle, (3) column-only shuffle, and (4) both row and column shuffle. Recall that unlike natural-language text, two dimension relational tables are permutation-invariant [table-gpt, cong2023observatory], meaning that shuffling rows and columns should generally not change their semantic meanings and the associated tasks (e.g., the tasks in Figure 1 will remain the same, even if the rows/columns in the associated tables are shuffled).
Table format variations. Figure 8 shows different models’ performance using common table input formats: Markdown, CSV, JSON, and HTML. We observe that unlike prior studies that showed notable sensitivity of LLMs [table-llm] to table formats, our results indicate that today’s frontier models, especially reasoning ones (GPT-5 and R1), are becoming less sensitive to these format variations (except HTML that still lags behind other formats). This reduces the need for format-specific optimization as models continue to improve. For chat models (GPT-5-Chat and DeepSeek-V3), we note that using JSON has an advantage on MMTU, mainly because it is easier for models to identify value/column correspondence in the JSON format, especially in long table context settings (e.g., in needle-in-a-haystack tasks like NIHT and NIHI shown in Table 1).
4.3 Error analysis
We observe that models are prone to errors when handling long table contexts, such as multiple tables or large tables with many rows and columns, as illustrated in Figure 6. Figure 9 (Right) shows a concrete example of the issue in the task of table-reshaping (TTBR in Table 1), where the reasoning trace shows that the model miscalculates the column index of the target column. This issue of long-context table is also notable in tasks like List-to-Table (L2T), Data Imputation (DI), Equi-Join (EJ), and NL-to-SQL (NS), etc.
A more detailed analysis of challenges in long-context tables is provided in Appendix E, where we examine a table-specific variant of the “needle-in-a-haystack” task, to understand why long-context tables remain challenging for models.
Figure 9: (Left): An example “table understanding” error in Data Imputation: while the model knows the person’s name for the missing cell, it misses the table context and uses an abbreviated form of the name, instead of the full-name in the ground-truth. (Right): Another example “table understanding” error in Table Reshaping: while the model knows the correct transforms to perform (“explode” in Pandas), it miscounts the column index for the target column “Season(s)”, which should be the sixth column instead of the fifth based on the reasoning trace, which leads to an incorrect transformation.
LLM-based error analysis. In addition, we complement the manual error analysis, by instructing LLM to inspect the question and ground-truth, to classify model prediction errors into these categories. We present the corresponding results in Appendix F in the interest of space.
5 Conclusion
Future work includes broadening the scope of table tasks to cover those that are important in practice but underrepresented in the existing research literature, as well as incorporating more subjective or creative tasks such as data generation, summarization, and enrichment. Another promising direction is the integration of multi-modal models for evaluation, which may have an advantage over text-only models on tasks that require two-dimensional table understanding.
Appendix A Broader Discussions
A.1 Open access to data and code
Our benchmark data is available at https://huggingface.co/MMTU-benchmark, and our evaluation code is available at https://github.com/MMTU-Benchmark.
A.2 License for existing assets
We provide comprehensive references to the sources of our data assets throughout this paper, with more details in [code-data]. We respect and conform to the licensing terms of existing datasets, whenever we can find such terms when we create our benchmark.
Appendix B Task overview
B.1 Column transformation
Program Transform by Example (PTBE) [ptbe-tde, ptbe-transform-prose, transform-column-devlin2017robustfill, transform-column-singh2016blinkfill, he2018transform]. This is a popular task studied in the data management and programming language literature. In this task, we are given a few example output values in a new output column (usually provided by users as demonstrations), and the task is to synthesize a transformation program using input tables, such that the produced output can match the user-given output examples. Figure 1 shows an example of this task. We use the benchmark datasets from [ptbe-tde, ptbe-transform-prose] for this task in MMTU.
Semantic Transform by example (STBE) [sj-sema-join, sj-dataxformer, semantic-join-dtt]. This task is similar to PFBE above, in that users also provide a few example output to demonstrate the intended transformations, but the target transformation requires “semantic” transformations (e.g., country to capital-city, or company to stock-ticker), that cannot be programmed using a syntactic program like in PFBE. We use the datasets from [sj-sema-join, sj-dataxformer]for our benchmark.
Formula by Context (FBC) [fbc-auto-formula, chen2021spreadsheetcoder]. This task is studied in the context of spreadsheets, where given a target spreadsheet cell in which users want to enter a spreadsheet formula, we are asked to predict the intended formula in the target cell. We use the 4 benchmark datasets in [fbc-auto-formula].
B.2 Table Transformation
Table Transform by Output Table (TTBT) [ttbt-autopandas, ttbt-scythe, transform-table-qbo, table-transform-sql-by-output]. In this task, we are given a pair of input/output tables, and the task is to infer a transformation program (in SQL or Pandas), which when executed on the input table, can generate the desired output table. Figure 1 shows an example of this task. We use the datasets from [ttbt-autopandas, ttbt-scythe].
Table Transform by Output Schema (TTBS) [tbps-auto-pipeline, sharma2023automatic, sharma2025datamorpher]. This task is similar to TTBT above, except that the output table is a schematic depiction of how the desired output table should look like, without being the exact output that the desired transformation program would generate on the given input (as it is sometimes hard to generate such output table without first producing the desired program). We use the dataset in [tbps-auto-pipeline]for this task.
Table Transform by Relationalization (TTBR) [ttbr-auto-tables, transform-table-barowy2015flashrelate, transform-table-foofah, xiong2025tablecanoniser, autoprep]. Because relational data analysis often requires input tables to be in a relational form, this task transforms data in non-relational semi-structured forms, into the standard relational form. The task is to predict such a relationalization transformation program, based on the characteristics of the input table. We use the dataset in [ttbr-auto-tables].
B.3 Table matching
Entity Matching (EM) [em-magellandata, em-deepmatcher, em-auto-em, em-ditto, em-book]. Entity matching is the task of determining whether rows or entities from two tables correspond to the same real-world entity. We use the datasets in  [em-magellandata, em-deepmatcher]for benchmarking.
Schema Matching (SM) [sm-valentine, sm-smurf, sm-other-1, schema-matching-cupid]. Schema matching tries to match relevant columns from two tables that map to the same semantic concept. Figure 1 shows an example of this task. We use datasets in [sm-valentine, sm-smurf]for MMTU.
Header Value Matching (HVM) [table-gpt, schema-mapping-survey]. In HVM, we are given a table without headers, and a shuffled list of the original headers in the table, where the task is to match column headers with column values in tables. We use the dataset in [table-gpt]for benchmarking.
B.4 Data cleaning
Data Imputation (DI) [webtables, tablib, table-gpt]. Data imputation is the intuitive task of predicting missing values in a relational table, based on the surrounding context of the table. Figure 1 shows an example of this task. We use the dataset from [webtables, tablib].
Error Detection (ED) [error-detection-survey, error-detect-uni-detect, error-detect-baran, error-detect-auto-test]. The task of Error detection aims to identify erroneous values in a table that are semantically inconsistent or anomalous with the rest of the column. We use the dataset in [error-detect-auto-test].
List to Tables (L2T) [l2t-tegra, list-extraction-2, list-extraction-google, list-other-1]. In List-to-table, the task is to segment records of data without clear separators, into columns, so that values in the same column are homogeneous, and the resulting table becomes a natural relational table, with values consistently aligned in homogenous columns. We use the dataset from [l2t-tegra].
B.5 Table join
Equal Join (EJ) [ej-auto-bi, ej-chen2014fast, ej-fk-ml, ej-hpi]. In Equal-join, we are given a collection of related tables, and the task is to identify all join relationships between the tables. Figure 1 shows an example of this task. We use the dataset in [ej-auto-bi].
Semantic Join (SJ) [sj-sema-join, sj-dataxformer, semantic-join-dtt, auto-map]. In Semantic Join, we are also tasked to join related tables together, but instead of the common equi-join, which is based on string-equality comparisons, the join relationship is based on semantic relatedness (e.g., country and capital city, or company and stock ticket). We use the same semantic-transformation datasets from [sj-sema-join, sj-dataxformer], but converting the input/output transformation columns, as join keys from two separate tables.
B.6 Column relationships
Arithmetic Relationship (AR) [explain-relationship, Auto-relate]. This task focuses on identifying arithmetic relationships from an underlying table. Figure 1 shows an example of this task.
String Relationship (SR) [Auto-relate, ptbe-tde, transform-col-flashfill]. This task focuses on identifying string transformation relationships from a given table.
Functional Relationship (FR) [Auto-relate, fd-1, fd-2]. This task focuses on identifying functional-relationships from input tables. We use the datasets from [Auto-relate] our as benchmarks for all three tasks.
B.7 Table understanding
Needle-in-a-haystack-table (NIHT). In this task, we create a variant of the popular “Needle-in-a-haystack” task in the context of tables, like described in detail in Appendix E. We use the dataset from [tablib]to construct tests in this task.
Needle-in-a-haystack-index (NIHI). In this task, we reverse the NIHT task above, and ask models to identify index positions corresponding to a given value in a table. We use the same dataset from [tablib]to construct tests for this task.
B.8 NL-2-Code
NL-2-SQL (NS) [ns-archer, ns-bird, ns-kaggledbqa, ns-spider, ns-wikisql]. NL-2-SQL is a popular task to translate natural-language questions into SQL statements that can execute given an input table. We use the benchmarks from [ns-archer, ns-bird, ns-kaggledbqa, ns-spider, ns-wikisql]in MMTU.
B.9 Table Question Answering (Table QA)
Table-QA (TQA) [tqa-tablebench, tqa-wikiqa, tqa-finqa, tqa-hitab]. Table QA is another popular task, where models are used to directly answer questions on a given input table, without code execution. We use benchmarks from [tqa-tabfact, tqa-tablebench, tqa-wikiqa, tqa-finqa]for this task.
Fact verification (FV) [tqa-tabfact, tablefact-2, tablefact-3]. Table fact verification is a variant of TQA, in which models are asked if a statement is refuted or supported by facts presented in an input table. We use data from [tqa-tabfact]for this task.
B.10 Knowledge-base mapping (KB Mapping)
Column type annotation (CTA) [semtab19, cta-autotype, cta-doduo, cta-sato, cta-sherlock]. This is a popular task where columns in a table is mapped to known entity types in a knowledge-base. We use the CTA dataset from [semtab19].
Column property annotation (CPA) [semtab19, cpa-2, cpa-3]. [semtab19]. The CPA task is to predict the relationship of a given pair of columns, based on known properties in a knowledge-base. We use the CPA dataset from [semtab19].
Cell entity annotation (CEA) [semtab19, turl, cea-1, cea-2, cea-3]. The CEA task predicts the knowledge-base entity id of a given cell in a table. We use the CEA dataset from [semtab19].
Appendix C Additional Statistics
Table 4 lists the exact ranges of token count calculated for MMTU questions, in each quartile and for each task.
@@ Appendix C Additional Statistics
Task Name Dataset Chat Model Reasoning Model
GPT-5-Chat DeepSeek-V3 GPT-5 Deepseek-R1
Data-transform-reshape Auto-Tables 0.500 0.282 0.710 0.126
Transform-by-output-target-schema commercial-pipelines 0.231 0.154 0.385 0.154
github-pipelines 0.253 0.224 0.269 0.213
Average 0.242 0.189 0.327 0.183
Transform-by-input-output-table AutoPandas 0.593 0.481 0.889 0.852
Scythe 0.667 0.600 0.950 0.883
Average 0.630 0.541 0.919 0.868
Entity-Matching Amazon-Google 0.872 0.856 0.938 0.868
BeerAdvo-RateBeer 0.966 0.955 0.989 0.943
DBLP-ACM 0.983 0.985 0.997 0.973
DBLP-Scholar 0.962 0.962 0.976 0.950
Fodors-Zagats 1.000 0.995 1.000 0.989
Walmart-Amazon 0.948 0.940 0.983 0.940
iTunes-Amazon 0.953 0.943 0.991 0.915
Average 0.955 0.948 0.982 0.940
header-value-matching TableGPT 0.699 0.633 0.818 0.678
Schema-Matching DeepMDatasets 1.000 1.000 1.000 1.000
HXD 0.928 0.945 0.937 0.937
Wikidata 0.897 0.882 0.912 0.931
assays 0.843 0.913 0.887 0.924
miller2 0.916 0.946 0.925 0.929
prospect 0.976 0.927 0.996 0.958
Average 0.927 0.935 0.943 0.947
Data-Imputation WebTable 0.493 0.483 0.642 0.499
tablib 0.704 0.627 0.843 0.664
Average 0.599 0.555 0.742 0.582
Error-Detect Relational-Tables 0.136 0.214 0.211 0.153
Spreadsheet-Tables 0.103 0.123 0.128 0.113
Average 0.119 0.168 0.170 0.133
List-to-table TEGRA 0.604 0.600 0.598 0.567
semantic-join DataXFormer 0.797 0.759 0.834 0.840
SEMA-join 0.899 0.889 0.947 0.928
Average 0.848 0.824 0.890 0.884
equi-join-detect Auto-BI 0.665 0.655 0.692 0.688
Data-transform-pbe TDE 0.487 0.419 0.708 0.614
Transformation-text 0.515 0.479 0.681 0.636
Average 0.501 0.449 0.694 0.625
Formula-prediction-context cisco-random 0.124 0.101 0.224 0.166
enron-random 0.051 0.043 0.163 0.105
pge-random 0.071 0.080 0.116 0.125
ti-random 0.076 0.051 0.179 0.140
Average 0.081 0.069 0.170 0.134
semantic-transform DataXFormer 0.423 0.419 0.507 0.443
SEMA-join 0.877 0.865 0.915 0.883
Average 0.650 0.642 0.711 0.663
Arithmetic-Relationship Auto-Relate 0.476 0.506 0.735 0.693
Functional-Dependency Auto-Relate 0.739 0.730 0.799 0.762
String-Relationship Auto-Relate 0.148 0.128 0.388 0.551
Table-needle-in-a-haystack MMTU 0.546 0.446 0.920 0.703
Table-Locate-by-Row-Col MMTU 0.570 0.517 0.941 0.710
NL2SQL Archer 0.327 0.183 0.394 0.404
KaggleDBQA 0.535 0.465 0.524 0.481
Spider 0.785 0.776 0.739 0.707
WikiSQL 0.745 0.733 0.732 0.704
bird 0.463 0.429 0.478 0.459
Average 0.571 0.517 0.574 0.551
Table-QA FinQA 0.211 0.229 0.497 0.403
TableBench 0.441 0.432 0.580 0.566
WikiQA 0.728 0.685 0.827 0.797
Average 0.460 0.449 0.635 0.589
Table-Fact-Verification TabFact 0.852 0.838 0.940 0.927
Column-type-annotation SemTab2019 0.554 0.409 0.462 0.371
Columns-property-anotation SemTab2019 0.296 0.258 0.492 0.280
Cell-entity-annotation SemTab2019 0.687 0.676 0.803 0.689